OpenStudy (anonymous):

Really Confused!!!!!!!!!!!!! G varies inversely with H, when H is 4 , G is .25 find the value of G when H is .3

OpenStudy (anonymous):

G varies inversely with H or we can say G * H= k where k is constant Now when H is 4 , G is .25 so 4*.25 =k or k=1 now we have GH= 1 so G = 1/H =1/0.3

OpenStudy (anonymous):

How do you know it will multiply

OpenStudy (anonymous):

G varies inversely with H, means when H increases G decreases...

OpenStudy (anonymous):

so G and H are inversely related...